# Combine and conquer: model averaging for out-of-distribution forecasting

### v1

@@ 3.5.1 Performance of individual sub-models
We see that for the DECISIONS data, DFT obtains a higher log-likelihood (LL) than NL, while MNL outperforms DFT in the LPMC data. In both studies, the two machine learning approaches obtain much higher LL than NL/MNL and DFT. Proportionally, the difference is larger in the DECISIONS case study than in the LPMC one, where in the former, MLP also has a bigger advantage over XGB. No formal statistical tests are carried out to compare the different model structures as these are not applicable to the machine learning structures.
@@
As a next step, Figure 3 looks at the performance of these models across the different distance segments obtained by splitting the data into deciles of the distance distribution. We do this separately for the estimation and validation data. A point to note when studying these results is that for the individual models, only segments 3 to 8 were used in estimating the model parameters, but the probabilities in segments 2 to 9 for the individual models also contributed to the model averaging, as explained in Figure 1. The values shown for segments 2 and 9 in the left two graphs in Figure 3 thus in effect also relate to out-of-distribution validation for the individual models.
Figure 3: Estimation and validation performance
@@
Starting with the graphs on the left, we see that, across the two datasets, the advantage for MLP and XGB that we noted at the overall level in Table 2 apply in each of the 6 distance segments used in estimation (i.e. segments 3-8). Across the distance segments, there are some where XGB outperforms MLP (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases. These findings already show that different models perform differently well across trip distances. Interestingly, while the LPMC results show that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
@@
We next turn to segments 2 and 9 in the left two graphs, giving us initial insights into out-of-distribution performance, remembering that the values shown here are the predicted probabilities for the chosen alternative, using the model parameters obtained when estimating the models on segments 3 to 8. In the DECISIONS data, we see all models perform very similarly in segment 2 (except for XGB), while in segment 9, NL clearly outperforms the data-driven models (MLP and XGB) and DFT. In the LPMC models, we see that MLP performs best in both segments 2 and 9, but the differences are very small.
@@
We next turn to the validation results, shown in the right two graphs in Figure 3. We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data. For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for the NL and DFT, with an average gain in per observation LL of $0.02752$ and $0.01811$, respectively, where this is most notable in segment 8. For MLP and XGB, we see small drops, with changes of $-0.0134$ and $-0.0058$, respectively. In the LPMC models, the changes are smaller, with gains in per observation LL of $0.0098$ for MNL, $0.01$ for DFT, $0.0043$ for MLP, and a drop by $-0.0008$ for XGB. Overall, these results raise no concerns about overfitting to the estimation data, at least within the distribution of distances covered by the estimation data.
@@
The findings for segments 2 and 9 on the validation data largely confirm those we saw in the two graphs on the left in Figure 3. We finally turn to segments 1 and 10. From a behavioural perspective, we see that for the DECISIONS data, mode choice on very short tips is much easier to predict. For short trips in the LPMC dataset, there is a clear advantage for the behavioural models. For the longest trips (segment 10), the trend is less clear. In the DECISIONS data, MLP performs comparatively well. In the LPMC data, both data-driven models do comparatively well, while MNL suffers a big drop.
@@ 3.5.2 Model averaging results
In model averaging, data covering segments 2-9 were used, but with model parameters for the individual models estimated on segments 3-8 only (cf. discussion around Figure 1). Table 3 shows the performance (in terms of LL on the estimation set) of the four individual models, alongside that of the model averaging (MA) structure. In both datasets, the MA structure clearly obtains a higher LL than any of the four individual models, although in LPMC, the difference to the best-fitting sub-model (MLP) is only around 0.13%. This in itself is simply a reflection of the overall dominance of MLP over the other three model types in segments 2-9 for this data, and further insights will be gained in model validation. A potential reason for this dominance could be the larger sample size compared to the DECISIONS data, which could enable MLP and XGB to better learn the data generating process.
@@
NL/MNL -3,397.316 (13.62%) -38,383.5 (3.91%)
DFT -3,451.558 (14.98%) -38,366.13 (3.86%)
MLP -3,019.805 (2.83%) -36,931.39 (0.13%)
XGB -3,202.91 (8.38%) -37,228.17 (0.92%)
MA -2,934.45 -36,884.39
The inclusion of segments 2 and 9 allows the model averaging structure to learn about out-of-distribution performance and hence guide the process of linking the weights given to individual models to the difference between the distance for that trip and the data used in estimation (i.e. segments 3 to 8). As explained in Section 3.4, a very flexible neural network was used to estimate this relationship, with no a priori assumptions about the functional form. The resulting relationship is illustrated in Figure 4.
For both datasets, we see that the MA process gives almost the entire weight to MLP and XGB for any trips with a distance contained within $D_{train,sub}$, i.e. distances on which the individual sub-models were estimated. In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP. These findings are entirely in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
@@
The final step in our analysis of results concerns the performance of the model averaging structure on the validation data. Table 4 shows the LL on the validation data, covering the entire range of distances, for the four sub-models and for model averaging. As in estimation, we see that the model averaging structure obtains the least negative LL. What is striking however is the contrast with the results on the estimation data, i.e. Table 2. With the exception of DFT on the LPMC data, the gains by model averaging over the sub-models are larger on the validation data than on the estimation data. This provides initial evidence to support the notion that model averaging is less susceptible to over-fitting on the estimation data than data-driven models.
Table 4: Log-Likelihood for model averaging: individual components and overall structure on validation data (covering distance segments 1-10), with gains by MA over individual models shown in brackets
@@
NL/MNL -1,948.29 (15.32%) -22,233.56 (14.1%)
DFT -1,980.411 (16.7%) -19,414.46 (1.62%)
MLP -1,722.19 (4.21%) -19,869.84 (3.88%)
XGB -1,895.97 (12.99%) -22,131.08 (13.7%)
MA -1,649.75 -19,099.30
There are two possible influences that could drive the differences between Table 3 and Table 4. The first of these is the potential that model averaging reduces the risk of over-fitting to the estimation data, which can be tested by comparing the performance on segments 2-9 between the estimation and validation data. The second is the potential that model averaging offers improved out-of-distribution performance, which can be tested by comparing the performance on segments 1 and 10 between the sub-models and the model averaging structure.
To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and validation. We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 7 out of 8 for the LPMC data. When we turn to the validation data, model averaging performs best (or joint best) in 9 out of 10 segments for the DECISIONS data, and 8 out of 10 segments for the LPMC data. These per segment statistics go hand in and hand with the best overall LL obtained by model averaging as reported in Table 3 and Table 4.
Figure 5: MA Estimation and validation performance
In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments where the per observation LL increases and segments where it decreases. No clear pattern emerges in the comparison across models and segments. Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.3419 vs -0.3628 for MLP). For the LPMC data, a similar story emerges. Model averaging is outperformed slightly by NL in segment 1, and slightly by DFT in segment 10, but overall performs the best of all models when looking at segments 1 and 10 together (-0.6078 vs -0.6153 for DFT).
Overall, these findings lead to the conclusion that there is little difference across the models in terms of out-of-sample performance (when looking at segments 2-9), aside from the potential suggestion that the risk of overfitting is reduced with more behavioural models, as already shown in Figure 3. There is no particular benefit for model averaging over other models in this context (given the dominance of the data-driven models), and nor was this the expectation. On the other hand, there is clear evidence that our model averaging approach offers benefits in terms of out-of-distribution performance, which was its intended aim.
@@ 4 Conclusions
This work has taken an important step forward in combining insights from different modelling approaches for travel demand forecasting. Specifically, we have shown that different models predict choices differently well depending on how far away from the estimation sample the prediction takes place. This result is not surprising in itself but is quantified by our work.
The main contribution comes in the development of an MLP-based model averaging approach that estimates weights for different models as a function of the distance away from the estimation data. Across two different case studies looking at mode choice, we see that data-driven machine learning approaches obtain almost the entire weight when making predictions for trips in areas similar to the trips in the estimation data. However, once we move outside that area and look at out-of-distribution prediction, the model averaging approach increases the weight to the models with behavioural underpinnings, sometimes substantially so.
In terms of prediction performance, we observe that model averaging offers the best overall performance on both the estimation and validation data. Crucially, this advantage is at its largest when making out-of-distribution predictions, in our case, predicting mode choice for trips whose distance is outside the range covered by the estimation data.
There are, of course, some opportunities for further refinement.
First, in our implementation, we used the same data to estimate the sub-models and the model averaging. In case of overfitting by one of the sub-models (an issue that did not arise in our case), this could create the risk that model averaging assigns disproportional weight to that model, which would jeopardise the benefits of the approach. In studies with larger sample sizes, a further splitting of the data could be considered.
Second, in applying our proposed model averaging approach, an analyst needs to define what distribution means. In our work, we have focussed on the obvious example of trip distance. A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes other than trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data. Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time.
Finally, we used MLP for the model averaging structure. Other meta-models, such as hybrid architectures approaches like TasteNetMNL (cf. Han et al., 2022), are an interesting avenue for future research.

### v2

@@ 3.5.1 Performance of individual sub-models
We see that for the DECISIONS data, NL obtains a higher log-likelihood (LL) than DFT, while DFT outperforms MNL in the LPMC data. In both studies, the two machine learning approaches obtain much higher LL than NL/MNL and DFT. Proportionally, the difference is larger in the DECISIONS case study than in the LPMC one. In the former, MLP outperforms XGB, while the reverse applies for the latter. No formal statistical tests are carried out to compare the different model structures as these are not applicable to the machine learning structures.
@@
As a next step, Figure 3 looks at the performance of these models across the different distance segments obtained by splitting the data into deciles of the distance distribution. We do this separately for the estimation and test data. A point to note when studying these results is that for the individual models, only segments 3 to 8 were used in estimating the model parameters, but the probabilities in segments 2 to 9 for the individual models also contributed to the model averaging, as explained in Figure 1. The values shown for segments 2 and 9 in the left two graphs in Figure 3 thus in effect also relate to out-of-distribution validation for the individual models.
Figure 3: Estimation and test performance
@@
Starting with the graphs on the left, we see that, across the two datasets, the advantage for MLP and XGB that we noted at the overall level in Table 3 apply in each of the 6 distance segments used in estimation (i.e. segments 3-8). Across the distance segments, there are some where XGB outperforms MLP in the DECISIONS data, or MLP outperforms XGB in LPMC (reversing the overall trend), while the overall trend between DFT and NL/MNL is also reversed in some cases. These findings already show that different models perform differently well across trip distances. Interestingly, while the LPMC results show that mode choice is easier to predict for longer trips (across segments 3-8), no clear trend emerges for the DECISIONS data.
@@
We next turn to segments 2 and 9 in the left two graphs, giving us initial insights into out-of-distribution performance, remembering that the values shown here are the predicted probabilities for the chosen alternative, using the model parameters obtained when estimating the models on segments 3 to 8. In the DECISIONS data, we see all models perform very similarly in segment 2, while in segment 9, DFT outperforms XGB. In the LPMC models, NL and DFT perform best in segment 2, while MLP does best in segment 9, but differences are again small.
@@
We next turn to the test results, shown in the right two graphs in Figure 3. We first focus on segments 3-8 with a view to understanding potential overfitting to the estimation data. For the DECISIONS data, we see some evidence of improved out-of-sample (vs within-sample) performance for all models, with an average gain in per observation LL of $0.096$ for NL, $0.095$ for DFT, $0.062$ for MLP, and $0.054$ for XGB. The same happens for LPMC models, the same is observed, with improvements by $0.12$ for NL and DFT, $0.11$ for MLP and $0.094$ for XGB. Overall, these results do not suggest any overfitting, at least within the distribution of distances covered by the estimation data.
@@
The findings for segments 2 and 9 on the test data largely confirm those we saw in the two graphs on the left in Figure 3, except for MLP and XGB performing more poorly in the test data for segment 9. We finally turn to segments 1 and 10. From a behavioural perspective, we see that for the DECISIONS data, mode choice on very short tips is much easier to predict. For short trips in the LPMC dataset, there is a clear advantage for the behavioural models. For the longest trips (segment 10), the trend is less clear. In the DECISIONS data, MLP performs comparatively well. In the LPMC data, both data-driven models do comparatively well, while MNL suffers a big drop.
@@ 3.5.2 Model averaging results
In model averaging, data covering segments 2-9 were used, but with model parameters for the individual models estimated on segments 3-8 only (cf. discussion around Figure 1). Table 4 shows the performance (in terms of LL on the estimation set) of the four individual models, alongside that of the model averaging (MA) structure. In both datasets, the MA structure clearly obtains a higher LL than any of the four individual models. It is noteworthy that this finding is robust even to variations in the performance of the individual models. In other words, the MA structure attains a similarly high log-likelihood even when one or more of the constituent models perform less well. This echoes the well-known principle in machine learning that an ensemble of weak learners can together form a strong learner. From a practical perspective, this suggests that analysts need not exhaustively optimise each individual model when developing a model-averaging framework.
@@
NL/MNL -3,134.33 (11.15%) -40,210.60 (4.49%)
DFT -3,061.83 (9.04%) -40,186.25 (4.43%)
MLP -2,846.66 (2.17%) -38,966.05 (1.44%)
XGB -2,999.89 (7.16%) -38,882.09 (1.23%)
MA -2,784.99 -38,404.69
The inclusion of segments 2 and 9 allows the model averaging structure to learn about out-of-distribution performance and hence guide the process of linking the weights given to individual models to the difference between the distance for that trip and the data used in estimation (i.e. segments 3 to 8). As explained in Section 3.4, a flexible neural network is used to estimate this relationship, with no a priori assumptions about the functional form. The resulting relationship is illustrated in Figure 4.
For both datasets, we see that the MA process gives almost the entire weight to MLP and XGB for any trips with a distance contained within $D_{train,sub}$, i.e. distances on which the individual sub-models were estimated. In both datasets, XGB gets higher weight for short trips within that part of the distance distribution compared to long trips, while for longer trips, almost the entire weight goes to MLP for DECISIONS, while it also increases for LPMC. These findings are in line with the pattern of prediction performance obtained for individual models, as shown on the left hand side in Figure 3.
@@
The final step in our analysis of results concerns the performance of the model averaging structure on the test data. Table 5 shows the LL on the test data, covering the entire range of distances, for the four sub-models and for model averaging. As in estimation, we see that the model averaging structure obtains the highest LL. What is notable is the contrast with the results on the estimation data, i.e. Table 3. For both datasets, the gains by model averaging over the best performing sub-model are larger on the test data than on the estimation data.
Table 5: Log-Likelihood for model averaging: individual components and overall structure on test data (covering distance segments 1-10), with gains by MA over individual models shown in brackets
@@
NL/MNL -825.88 (6.12%) -11,764.56 (5.30%)
DFT -805.78 (3.78%) -11,426.72 (2.50%)
MLP -804.91 (3.68%) -11,448.70 (2.68%)
XGB -829.64 (6.55%) -12,082.56 (7.79%)
MA -775.32 -11,141.55
There are two possible influences that could drive the differences between Table 4 and Table 5. The first of these is the potential that model averaging reduces the risk of over-fitting to the estimation data, which can be tested by comparing the performance on segments 2-9 between the estimation and test data. The second is the potential that model averaging offers improved out-of-distribution performance, which can be tested by comparing the performance on segments 1 and 10 between the sub-models and the model averaging structure.
To support this discussion, Figure 5 shows the per observation log-likelihoods in the different segments for estimation and test. We first see that, in estimation, model averaging outperforms (or equals) the four sub-models in 4 out of 8 estimation segments for the DECISIONS data, and 6 out of 8 for the LPMC data. When we turn to the test data, model averaging performs best (or joint best) in 2 out of 10 segments for the DECISIONS data, and 5 out of 10 segments for the LPMC data. These per segment statistics go hand in hand with the best overall LL obtained by model averaging as reported in Table 4 and Table 5.
Figure 5: MA Estimation and test performance
We earlier raised two possible reasons for why the advantages of MA might be larger on the test data than on the validation data. In relation to over-fitting to the estimation data, we see that, for all four sub-models as well as for the model averaging structure, there are segments (out of 2-9) where the per observation LL increases and segments where it decreases. No clear pattern emerges in the comparison across models and segments. Turning to the second potential reason, namely out-of-distribution performance, we observe that for the DECISIONS data, model averaging is outperformed by NL in segment 1, but obtains by far the best performance in segment 10, where its combined performance on segments 1 and 10 is much higher than for any of the other models (-0.2582 vs -0.2944 for XGB). For the LPMC data, a similar story emerges. Model averaging is outperformed slightly by NL in segment 1, but obtains by far the best performance in segment 10. Again, overall, MA performs the best of all models when looking at segments 1 and 10 together (-0.5793 vs -0.6025 for DFT).
Overall, these findings lead to the conclusion that there is little difference across the models in terms of out-of-sample performance when looking at segments 2-9. On the other hand, there is clear evidence that our model averaging approach offers benefits in terms of out-of-distribution performance, which was its intended aim.
@@ 4 Conclusions
This work has taken an important step forward in combining insights from different modelling approaches for travel demand forecasting. Specifically, we have shown that different models predict choices differently well whether we are within the ranges of the estimation data, or outside. This result is not surprising in itself but is quantified by our work.
The main contribution comes in the development of an MLP-based model averaging approach that estimates weights for different models as a function of the distance away from the estimation data. Across two different case studies looking at mode choice, we see that data-driven machine learning approaches obtain almost the entire weight when making predictions for trips in areas similar to the trips in the estimation data. However, once we move outside that area and look at out-of-distribution prediction, the model averaging approach increases the weight for the models with behavioural underpinnings, sometimes substantially so.
In terms of prediction performance, we observe that model averaging offers the best overall performance on both the estimation and test data. Crucially, this advantage is at its largest when making out-of-distribution predictions, in our case, predicting mode choice for trips whose distance is outside the range covered by the estimation data.
There are, of course, some opportunities for further refinement. The key one relates to the fact that, in applying our proposed model averaging approach, an analyst needs to define what distribution means. In our work, we have focused on the obvious example of trip distance, for which there were pragmatic reasons, but also given that trip distance embodies key information in relation to similarity between trips. A core opportunity for future work is thus the investigation of out-of-distribution performance considering attributes other than trip distance alone, for example, going outside the range of socio-economic characteristics covered by the estimation data, or combinations of socio-economic characteristics not covered in the estimation data. Out-of-distribution could also take on a multivariate nature, thus covering multiple trip characteristics at the same time. This, however, then also necessitates care by the analyst in defining what constitutes within-distribution and out-of-distribution.
Finally, we employed an MLP as the meta-model within the model averaging framework. The literature on ensemble and stacking architectures in machine learning offers a broad range of methodological extensions that could further enhance such meta-modelling approaches (e.g. Shaikh et al., 2024). Future research could, for instance, investigate deep stacking networks, Bayesian ensemble formulations, or ensemble distillation methods to improve model robustness and generalisation, particularly in data-scarce settings. Furthermore, in cases where a larger number of explanatory variables are available and one wishes to understand their importance, more structured meta-models, such as a logistic regression model or hybrid models, such as TasteNet Han et al., 2022 can be explored.